### v1

# Soft Token Attacks Cannot Reliably Audit Unlearning in Large Language Models

Abstract
Large language models (LLMs) have become increasingly popular. Their emergent capabilities can be attributed to their massive training datasets. However, these datasets often contain undesirable or inappropriate content, e.g., harmful texts, personal information, and copyrighted material. This has promoted research into machine unlearning that aims to remove information from trained models. In particular, approximate unlearning seeks to achieve information removal by strategically editing the model rather than complete model retraining.
Recent work has shown that soft token attacks (STA) can successfully extract purportedly unlearned information from LLMs, thereby exposing limitations in current unlearning methodologies. In this work, we reveal that STAs are an inadequate tool for auditing unlearning. Through systematic evaluation on common unlearning benchmarks (Who Is Harry Potter? and TOFU), we demonstrate that such attacks can elicit any information from the LLM, regardless of (1) the deployed unlearning algorithm, and (2) whether the queried content was originally present in the training corpus. Furthermore, we show that STA with just a few soft tokens ($1-10$) can elicit random strings over 400-characters long. Thus showing that STAs are too powerful, and misrepresent the effectiveness of the unlearning methods.
Our work highlights the need for better evaluation baselines, and more appropriate auditing tools for assessing the effectiveness of unlearning in LLMs.
Soft Token Attacks Cannot Reliably Audit Unlearning in Large Language Models
Haokun Chenthanks: Work done during an internship at Intel Labs. LMU Munich haokun.chen@campus.lmu.de                        Sebastian Szyller Intel Labs contact@sebszyller.com
Weilin Xu Intel Labs weilin.xu@intel.com                        Nageen Himayat Intel Labs nageen.himayat@intel.com
1 Introduction
In recent years, large language models (LLMs) have undergone substantial advancements, leading to enhanced performance and widespread adoption. LLMs have demonstrated exceptional performance in various downstream tasks, such as machine translation Zhu et al. (2023), content generation Acharya et al. (2023), and complex problem-solving Xi et al. (2025). Their performance is attributed to their large-scale architectures that require datasets consisting of up to billions of tokens to train effectively Kaplan et al. (2020). These datasets are typically derived from large-scale corpora sourced from public internet text. However, such datasets inadvertently contain harmful or inappropriate content, including instructions for hazardous activities (e.g., bomb-making), violent or explicit material, private information, or copyrighted content unsuitable for applications. Given the sensitive nature of such data, it may be necessary to remove it from the LLM to comply with the local regulations, or internal company policies.
Machine unlearning is a tool for removing information from models Cao and Yang (2015); Bourtoule et al. (2021a). Approximate unlearning usually refers to removing information from models without resorting to retraining them from scratch Zhang et al. (2024a); Eldan and Russinovich (2023a); Izzo et al. (2021), ensuring that the resulting model deviates from a fully retrained version within a bounded error. While numerous studies have proposed various unlearning algorithms, most lack formal guarantees of effectiveness. In fact, prior research has demonstrated that many unlearning techniques can be circumvented through simple rephrasings of the original data Shi et al. (2024). Recent work has shown that a soft token attack (STA) can be used to elicit harmful completions and extract supposedly unlearned information from models Schwinn et al. (2024); Zou et al. (2024).
In this work, we introduce a simple framework for auditing unlearning and demonstrate that STAs are overly powerful, and hence, inappropriate for verifying the effectiveness of approximate unlearning. We show that the auditor can elicit any information from the model, regardless of its training data. Our work highlights the need for better unlearning auditing baselines and methodologies.
We claim the following contributions:
  1. 1.
    We introduce a simple auditing framework for unlearning in LLMs (Section 3.2).
  2. 2.
    We show that STAs effectively elicit unlearned information in all tested unlearning methods and benchmark datasets (Who Is Harry Potter?, and TOFU). Additionally, we show that STAs also elicit information in the base models that were not fine-tuned on the benchmark datasets (Section 4.2).
  3. 3.
    We further demonstrate that the STAs are inappropriate for evaluating unlearning – we show that a single soft token can elicit $150$ random tokens, and ten soft tokens can elicit over $400$ random tokens (Section 4.3).
The remainder of this paper is organized as follows: In Section 2 we provide an overview of the necessary background, and the related work. Section 3 introduces a general auditing framework, and instantiates it using STA. In Section 4, we demonstrate the efficacy of STA, and subsequently its failure, as a tool for auditing unlearning in LLMs. In Section 5 we discuss additional considerations for auditing unlearning in LLMs. We conclude the paper in Section 6, and highlight some limitations in Section 7.
2 Background & Related work
2.1 Background
Figure 1: Overview of the auditing process using $A_{\textsl{STA}}$. For a perfect unlearning method, $A_{o}$ always correctly audits the model. On the other hand, $A_{\textsl{STA}}$ can elicit the completion regardless of the information in the model – the audit is ineffective.
Large language models (LLMs) process input text through an auto-regressive framework. Given an input sequence of tokens $x_{1:t}$, the model computes the conditional probability distribution $p(x_{t+1}|x_{1:t})$ over the vocabulary at each time-step. The likelihood of the sequence is given by:
$\log p(x_{t+1}|x_{1:t})=\sum_{t=1}^{T}\log p(x_{t}|x_{1:t-1})$ (1)
At inference time, the tokens is generated iteratively by sampling the next token $x_{t+1}$ from this distribution (e.g., via greedy decoding or nucleus sampling Holtzman et al. (2019)), then appending it to the context $x_{1:t}$ for the subsequent step.
Machine unlearning is a tool for removing information from models. Consider a machine learning model $f$ optimized over a training dataset $D_{train}$. When a data owner submits an unlearning request to remove a specified subset $D_{forget}\in D_{train}$, the objective of machine unlearning is to produce an unlearned model $f_{u}$ that eliminates the influence of $D_{forget}$. Machine unlearning methodologies are categorized into two paradigms – exact, and approximate unlearning.
Exact unlearning ensures the output distribution of $f_{u}$ is statistically indistinguishable from that of a model retrained exclusively on the retained dataset $D_{retain}=D_{train}/D_{forget}$. This guarantees provable data removal, satisfying:
$p(f_{u}(x)=y)$ $=p(f_{ret}(x)=y)\quad$ (2)
$s.t.\quad\forall(x,y)\in D_{train},$
where $f_{ret}$ denotes a model retrained from scratch on $D_{retain}$ – which is the most straightforward way of achieving exact unlearning.
The process can be made more efficient by splitting the $D_{train}$ into overlapping chunks, and training an ensemble of models Bourtoule et al. (2021b). During an unlearning request, only the models containing the requested records need to be retrained. For certain classes of models, it is possible to achieve exact unlearning without retraining, e.g. ECO adapts the Cauwenberghs and Poggio (CP) algorithm for exact unlearning within LeNet Huang et al. , and MUSE relabels the target data to achieve unlearning for over-parameterized linear models Yang et al. (2024).
Approximate unlearning, sometimes called in-exact unlearning, relaxes the strict equivalence requirement, instead only requiring that $f_{u}$ approximates $f_{ret}$ within some bounded error. This paradigm relies on empirical metrics or probabilistic frameworks. In LLMs, approximate unlearning is typically accomplished by overwriting the information in the model Eldan and Russinovich (2023a); Wang et al. (2024), guiding the model away from it Feng et al. (2024), or editing the weights and/or activations Liu et al. (2024); Bhaila et al. (2024); Li et al. (2024); Tamirisa et al. (2024); Huu-Tien et al. (2024); Ashuach et al. (2024); Meng et al. (2022a, b).
2.2 Related work
While advances have been made in developing machine unlearning algorithms for LLMs, rigorous methodologies for auditing the efficacy of unlearning remain understudied. Adversarial soft token attacks (STAs) Schwinn et al. (2024) and 5-shot in-context prompting Doshi and Stickland (2024) have been shown to recover unlearned knowledge in models. When model weights can be modified, techniques such as model quantization Zhang et al. (2024e) and retraining on a partially unlearned dataset  Łucki et al. (2024); Hu et al. (2024) have also proven effective in recalling forgotten information.  Lynch et al. (2024) examined eight methods for evaluating LLM unlearning techniques and found that their latent representations remained similar. News and book datasets are used to analyze unlearning algorithms from six different perspectives Shi et al. (2024). It was shown that fine-tuning on unrelated data can restore information unlearned from the LLM Qi et al. (2024), indicating the existing unlearning methods do not actually remove the information but learn a refusal filter instead. Several benchmarks have been developed to evaluate the existing unlearning algorithms. Besides, an unlearning benchmark was introduced based on fictitious author information Maini et al. (2024a). For real-world knowledge unlearning, Real-World Knowledge Unlearning (RWKU) used 200 famous people as unlearning targets Jin et al. (2024), while WDMP focused on unlearning hazardous knowledge in biosecurity, cybersecurity Li et al. (2024).
3 Auditing with Soft Token Attacks
3.1 Adversarial prompts
An adversarial prompt $x_{a}$, is an input prompt to the LLM, obtained by applying the transform $T(\cdot)$ to the base prompt $x_{p}$, $x_{a}=T(x_{p},\textsl{aux})$ in order to elicit a desired completion $c$. $T$ can be any function that swaps, removes or adds tokens; aux denotes any additional needed information. However, such arbitrary attacks are expensive to optimize [^1] Suffix-only attacks allow efficient use of the KV-cache., and difficult to reason about. In practice, $T$ optimizes an adversarial suffix $x_{s}$ that is appended to $x_{p}$ to elicit $c$ Zou et al. (2023). Specifically, we optimize the probability:
$Prob=P(c|x_{p}\oplus x_{s}).$ (3)
An adversary with white-box access to the LLM, can instead mount the attack in the embedding space i.e. modify the soft tokens:
$Prob=P(c|embed(x_{p})\oplus embed(x_{s})).$ (4)
In this case, $T$ uses the gradient from the LLM to update $x_{s}$.
3.2 Unlearning auditor
An oracle auditor $A_{o}$ takes an unlearned model $f_{u}$ and the candidate sentences $x_{c}\in X_{c}$ and outputs a ground truth, binary decision $a=\{0,1\}$ indicating whether the given records was part of $D_{train}$ of:
$a=A_{o}(f_{u},X_{c}=D_{forget},\textsl{aux})$ (5)
$A_{o}$ is unrealistic in many scenarios; however, it can be easily instantiated for exact unlearning where $A_{o}$ knows the training data associated with $f$: $aux=\{D_{retain}\}$.
On the other hand, a realistic unlearning $A_{u}$ takes an $f_{u}$, and $D_{forget}$ and outputs a score $s=(0,1)$ indicating whether the records were in $D_{train}$:
$s=A_{u}(f_{u},\{x_{c}\}=D_{forget},\textsl{aux}=\emptyset)$ (6)
$A_{u}$ represents cases where users remove information from models that they did not create, e.g. to prevent harmful outputs.
In this work, we instantiate the soft token attack auditor $A_{\textsl{STA}}$ based on the soft token attacks (STAs) against unlearning Schwinn et al. (2024); Zou et al. (2024). Our auditor compares the relative difficulty of eliciting $c$ for $f_{ft}$ and $f_{u}$. The unlearning procedure is effective if eliciting completions using $f_{u}$ is more difficult than $f_{ft}$.
$s=A_{\textsl{STA}}(f_{u},\{x_{c}\}=D_{forget},\textsl{aux}=\{f_{ft}\}).$ (7)
Figure 1 gives a complete overview of the auditing procedure, and the difference between $A_{0}$ and $A_{\textsl{STA}}$. In Table 1, we summarize the notation.
In the next Section, we show that $A_{\textsl{STA}}$ cannot reliably audit LLMs.
4 Evaluation
4.1 Experiment setup
Datasets. To attack unlearning and evaluate its effectiveness, we use two popular benchmark datasets: 1) Who Is Harry Potter?Eldan and Russinovich (2023a), a benchmark that intends to remove information about the world of Harry Potter and the associated characters; WHP hereafter. 2) TOFUMaini et al. (2024a), a dataset of fictional writers that are guaranteed to be absent in the LLM’s training data [^2] This cannot be guaranteed for models published after TOFU..
WHP does not publish a complete dataset. For that reason we use the passages included in the associated Hugging Face page Eldan and Russinovich (2023b). Additionally, we augment it with $20$ $(x_{p}\rightarrow c)$ pairs generated with Llama2-7b-chat-hf. These contain general trivia about the Harry Potter universe.
For TOFU, we use the $10\%$ forget to $90\%$ retain split provided by the authors Maini et al. (2024b).
Models & environment. For all experiments, we use Llama-2-7b-chat-hf Touvron et al. (2023) (Llama2), and Llama-3-8b-instruct Meta (2024) (Llama3) downloaded from Hugging Face. We get the unlearned WHP model from it’s Hugging Face repository Eldan and Russinovich (2023b) (Llama2-WHP).
We implement STA using LLMart Cornelius et al. (2025) – a PyTorch and Hugging Face-based library for crafting adversarial prompts. We use implementations of the unlearning methods from the TOFU Maini et al. (2024c), and NPO Zhang et al. (2024b) repositories. We benchmark the attack against seven different unlearning algorithms: gradient ascent (GA), gradient difference (GDF) Liu et al. (2022), refusal (IDK) Rafailov et al. (2024), knowledge distillation (KL) Hinton (2015), negative preference optimization (NPO) Zhang et al. (2024c), NPO-GDF, NPO-KL.
We run our experiments on a machine equipped with Intel Xeon Gold 5218 CPU, eight NVIDIA A6000, and 256 GB of RAM.
STA soft token attack
$A_{o}$ oracle auditor
$A_{\textsl{STA}}$ STA auditor
$x_{p}$ base prompt (benign)
$x_{s}$ adversarial suffix
$x_{a}$ adversarial prompt ($x_{p}\oplus x_{s}$)
$c$ target completion
$f_{\emptyset}$ base model
$f_{ft}$ fine-tuned model
$f_{u}$ unlearned model
$f_{u-*}$ model unlearned using *
$D_{train}$ training data
$D_{forget}$ forget data
$D_{retain}$ retain data
Table 1: Summary of the notation. ’*’ is replaced with the specific unlearning method.
4.2 Auditing with attacks
Who Is Harry Potter?. To elicit information about the Harry Potter universe, we initialize the soft tokens using randomly selected hard tokens, and append them to the prompt ($embed(x_{a}=x_{p}\oplus x_{s})$). We then train the soft prompt using AdamW Loshchilov and Hutter (2019) for up to $3000$ iterations; using learning $0.005$, and default $\beta{s}$ – we reiterate that the $x_{p}$ does not change, only the embedded suffix does. If the optimization fails, we double the the number of soft tokens up to the maximum of $16$. We rerun the experiment five times, and report the mean and standard deviation across all prompt and reruns. In Table 2 we report the average number of soft tokens needed to elicit a completion. WHP * denotes the unlearned model with different prompt templates.
Our results show that all information can be generated with $\approx 4-6$ added soft tokens. For all pairs of models, we conduct a t-test under the null hypothesis $\mathcal{H}_{0}$ of equivalent population distributions with $\alpha=0.05$. We use an unpaired Welch’s t-test since sample variances are not equal WELCH (1947). We cannot reject the hypothesis for any of the pairs i.e. $p>0.05$. In other words, for all models, there is not enough evidence to say that eliciting completions is more difficult.
[…]
Prompt Model
template Llama2-WHP Llama3
N/A N/A $5.61\pm 6.32$
WHP $4.63\pm 3.69$ N/A
[…]
WHP +chat $4.12\pm 5.53$ N/A
Table 2: Number of soft tokens required to elicit a completion for a fixed number of iterations. Soft tokens are appended to the prompt. Results are averaged over all prompts in the WHP set and over five runs for each prompt. When we increase the maximum iterations to $10,000$ we can elicit all completions with $1-2$ soft tokens. We do not report the results for Llama2 hf because it was used to generated the data. For comparison we also report the results for Llama3 without any prompt template (N/A).
Unlearning method Model
Llama2 Llama3
$f_{\emptyset}$ (none) $3.07\pm 3.25$ $3.11\pm 3.15$
$f_{ft}$ (none) $2.95\pm 3.35$ $3.21\pm 3.19$
$f_{u-IDK}$ $3.40\pm 3.20$ $3.33\pm 3.09$
$f_{u-GA}$ $3.34\pm 3.97$ $3.21\pm 3.87$
$f_{u-GDF}$ $3.06\pm 3.34$ $3.11\pm 3.40$
$f_{u-KL}$ $3.08\pm 3.31$ $3.12\pm 3.17$
$f_{u-NPO}$ $3.11\pm 3.27$ $3.12\pm 3.27$
$f_{u-NPO-GDF}$ $3.15\pm 3.24$ $3.16\pm 3.16$
$f_{u-NPO-KL}$ $3.23\pm 3.62$ $3.24\pm 3.57$
Table 3: Number of soft tokens required to elicit a completion for a fixed number of iterations. Soft tokens are appended to the prompt. Results are averaged over all prompts in the TOFU set and over five runs for each prompt. When we increase the maximum iterations to $10,000$ we can elicit all completions with $1-2$ soft tokens.
TOFU. For TOFU, we follow the same setup as for WHP– we initialize soft tokens using random hard tokens, and append them; we then train the soft prompt using AdamW for up to $3000$ iterations; using learning $0.005$, and $\beta{s}=(0.9,0.999)$; we double the soft tokens if the optimization fails; we rerun the experiments five times and report the averages across all prompts. In Table 3, we report the number of soft tokens required to elicit the completions. $f_{\emptyset}$ refers to the unmodified baseline model, $f_{ft}$ corresponds to the models fine-tuned on TOFU, followed by the unlearned models.
For all methods, we can elicit the completions with $\approx 3$ appended soft tokens. Similarly to WHP, for all possible pairs of models (within the same architecture), we conduct a t-test under the null hypothesis $\mathcal{H}_{0}$ of equivalent population distributions with $\alpha=0.05$. We use an unpaired Welch’s t-test since sample variances are not equal.
We cannot reject the hypothesis for any of the pairs i.e. $p>0.05$; $f_{u-IDK}$ vs $f_{ft}$ (for Llama2) gives the lowest p-value of $0.509$. In other words, for all models (regardless if trained on TOFU, or used unlearning method), there is not enough evidence to say that eliciting completions is more difficult.
One could argue that used unlearning methods are not effective (when comparing $f_{ft}$ vs $f_{u-*}$), hence they require similar numbers of soft tokens. [^3] Most of these approaches have been in fact been shown ineffective, and susceptible to simple paraphrasing. However, the same holds when compared to $f_{\emptyset}$. In the next section, we demonstrate that the result cannot be attributed to the (in-)effectiveness of the unlearning methods but rather the power of STA.
4.3 Eliciting random strings
To further show the power of STAs, we use them to elicit random strings. Unlike natural text, the chance of a random string appearing in the training set is negligible. Also, preceding tokens do not inform the selection of the next token. In the following experiments, we pick characters uniformly at random in the range 33-126 of the ASCII table asciitable.com .
To elicit a random string, we initialize the soft prompt using randomly selected tokens. Unlike in the WHP and TOFU experiments, there is only the soft prompt. We then train the soft prompt using AdamW for up to $3000$ iterations per soft token; using learning rate $0.005$, and $\beta{s}=(0.9,0.999)$.
In Figure 2, we report the longest elicited string for a given number of soft tokens. We repeat the experiment five times – e.g., the first marker implies that for each of the five tested random strings of length $150$, we found an effective soft prompt. We observe that not all initializations and seed configurations succeed, in which case a run needs to be restarted with a different seed. If the loss plateaus around $25\%$ of the iterations, we restart the run. However, no single string was restarted more than ten times. We experimented with learning rate schedulers but they did not improve the search.
Our results show that STAs can be used to elicit completely random strings, thus undermining their application for auditing unlearning. Due to limited computational resources and long run-times, our evaluation is limited to $10$ soft tokens, and $400$-character long random strings. This does not provide a bound on the longest string that can be elicited.
Next, we aim to answer why eliciting strings is possible. Prompt-tuning Lester et al. (2021) is a performance efficient fine-tuning technique in which instead of training all weights, one trains only a soft prompt added to the input. STAs can be viewed as an extreme case of prompt-tuning, where instead of training over many prompts, one trains an attack per each prompt. Hence, an LLM that outputs a completion that it was trained on is an expected behavior. We urge against misinterpreting the results and declaring techniques ineffective.
Figure 2: Numbers of random characters generated for the given soft prompt length. A single soft token can force over $150$ random characters – more than any text in the TOFU benchmark. With $10$ soft tokens, it is possible to generate over $400$ random characters.
5 Discussion & Conclusion
Auditing with hard prompts. Attacks such as greedy coordinate gradient Zou et al. (2023) optimize the attack prompt in the hard token space instead of the soft token space. Hence, they are weaker at eliciting completions. On one hand, this might make them more suitable for auditing unlearning. On the other hand, due to their computational requirements, they are often used to force only the beginning of a harmful completion (e.g. Sure, here’s how to build a bomb…) with the hope that the LLM follows. It is unclear whether this would be sufficient to produce specific unlearned passages. We see it as an interesting direction for future work.
Unlearning vs jail-breaking. Our findings are applicable to the jail-breaking community as well. Prior work Zhang et al. (2024d) hinted that unlearning and preventing harmful outputs can be viewed as the same task – removing or suppressing particular information. STAs and fine-tuning attacks Hu et al. (2024) are useful tools for evaluating LLMs in powerful threat models. It was shown that fine-tuning on benign data, or data unrelated to the unlearning records (for jail-breaking and unlearning respectively) can restore undesirable behavior Łucki et al. (2024).
Variation in gradient-based learning. Prior work showed that removing training records from the training set, and repeating the training can result in the same final model Thudi et al. (2022) depending on the random seed. Even though a record was part of the training run, its influence might be minimal, making unlearning unnecessary. Similarly, it was shown that SGD has intrinsic privacy guarantees, assuming there exists a group of similar records Hyland and Tople (2022). Thus, algorithmic auditing of unlearning might not be possible, and one would have to rely on verified or attested procedures instead Eisenhofer et al. (2023), regardless of their impact on the model.
Distinguishing learned soft tokens. Even though, in most our results, the number of soft tokens required to elicit a completion is the same, we attempted to distinguish between them. To this end, we take all single-token STAs optimized for TOFU (Table 3) and assign a label $y=\{0,1\}$: $y=0$ for $f_{\emptyset}$, and $y=1$ for $f_{ft}$ and the unlearned models. We then train a binary classifier using $f_{\emptyset}$ and $f_{ft}$. While we are able to overfit it and distinguish between $f_{\emptyset}$ and $f_{ft}$, we were not able to train a model that would generalize to the unlearned models, and decisively assign a class. Our approach is similar to Dataset Inference Maini et al. (2021, 2024d) which showed there can be distributional differences between the models, depending on the data they were trained on. Further investigation into what soft tokens are learned during the audit is an interesting direction for future work.
6 Conclusion
In this work, we show that soft token attacks (STAs) cannot reliably distinguish between base, fine-tuned, and unlearned models. In all cases, the auditor can elicit all unlearned information by appending optimized soft prompts to the base prompt. Additionally, we show that STA with a single soft token can elicit $150$ random characters, and over $400$ with soft tokens.
Our work demonstrates that machine unlearning in LLMs needs better evaluation frameworks. While many unlearning methods can be broken by simple paraphrasing of original prompts, or by fine-tuning on partial unlearned data or even unrelated data, STA misrepresents their efficacy.
7 Limitations & ethical considerations
Limitations. Due to computational constraints our work is limited to 7-8 billion parameter models. Nevertheless, given that LLMs’ expressive power increases with size Kaplan et al. (2020), our results should hold for larger LLMs. Our evaluation with random strings could be extended to verify if there is a clear and generalizable dependency between the number of soft tokens and the maximum number of generated characters.
Ethical considerations. In this work, we show that an auditor (a user) with white-box access to the model, and sufficient compute can elicit any text from the LLM. While it does require knowing the target completion for a given prompt, it is likely that partial completions might be enough, thus allowing the user to elicit harmful information. This may be particularly dangerous in settings where the user has approximate knowledge of the information that had been scrubbed off the LLM.
8 Acknowledgments
We thank Cory Cornelius and Marius Arvinte for valuable discussions on this work.