### v2

# Soft Token Attacks Cannot Reliably Audit Unlearning in Large Language Models

Abstract
Large language models (LLMs) are trained using massive datasets. However, these datasets often contain undesirable content, e.g., harmful texts, personal information, and copyrighted material. To address this, machine unlearning aims to remove information from trained models. Recent work has shown that soft token attacks (STA) can successfully extract unlearned information from LLMs. In this work, we show that STAs can be an inadequate tool for auditing unlearning. Using common unlearning benchmarks, i.e., Who Is Harry Potter? and TOFU, we demonstrate that, in a strong auditor setting, such attacks can elicit any information from the LLM, regardless of (1) the deployed unlearning algorithm, and (2) whether the queried content was originally present in the training corpus. Also, we show that STA with just a few soft tokens (1101-10) can elicit random strings over 400-characters long. Thus showing that STAs must be used carefully to effectively audit unlearning. Example code can be found at https://github.com/IntelLabs/LLMart/tree/main
/examples/unlearning
Soft Token Attacks Cannot Reliably Audit Unlearning in Large Language Models
Haokun Chenthanks: Work done while at Intel Labs. Ludwig Maximilian University of Munich Munich Center for Machine Learning (MCML) haokun.chen@campus.lmu.de          Sebastian Szyller11footnotemark: 1 Aalto University contact@sebszyller.com
Weilin Xu Intel Labs weilin.xu@intel.com          Nageen Himayat Intel Labs nageen.himayat@intel.com
1 Introduction
Large language models (LLMs) excel in many downstream tasks, e.g., machine translation Zhu et al. (2023), content generation Acharya et al. (2023), and complex problem-solving Chen et al. (2024). Their performance is attributed to their large-scale architectures that require datasets consisting of up to trillions of tokens to train effectively Kaplan et al. (2020). These datasets are typically derived from large-scale corpora sourced from public internet text. However, such datasets can contain undesirable content, e.g., instructions for building weapons, violent or explicit material, private information, or copyrighted content. Given the sensitive nature of such data, it may be necessary to remove it from the LLM to comply with the local regulations, or internal company policies.
Machine unlearning is a tool for removing information from models Cao and Yang (2015); Bourtoule et al. (2021a). Approximate unlearning usually refers to removing information from models without retraining them from scratch Zhang et al. (2024a); Eldan and Russinovich (2023a); Izzo et al. (2021), ensuring that the resulting model deviates from a fully retrained version within a bounded error. While numerous studies have proposed various unlearning algorithms, most lack formal guarantees. Prior research has demonstrated that many unlearning techniques can be circumvented through rephrasing of the original data Shi et al. (2024). Recent work has shown that a soft token attack (STA) can be used to elicit harmful completions and extract supposedly unlearned information from models Schwinn et al. (2024); Zou et al. (2024).
In this work, we introduce a simple framework for auditing unlearning, and demonstrate that STAs are inappropriate for verifying the effectiveness of approximate unlearning in a strong auditor setting. We show that the auditor can elicit any information from the model, regardless of its training data. We claim the following contributions:
  1. 1.
    We show that STAs effectively elicit unlearned information in all tested unlearning methods and benchmark datasets (Who Is Harry Potter?, and TOFU). Additionally, we show that STAs also elicit information in the base models that were not fine-tuned on the benchmark datasets (Section 5.2).
  2. 2.
    We further demonstrate that the STAs are inappropriate for evaluating unlearning – we show that a single soft token can elicit $150$ random tokens, and ten soft tokens can elicit over $400$ random tokens (Section 5.3).
2 Background
Adversarial prompt $x_{a}$ is an input prompt to the LLM, obtained by applying the transform $T(\cdot)$ to the base prompt $x_{p}$: $x_{a}=T(x_{p},\textsl{aux})$ to elicit a desired completion $c$. $T$ can be any function that swaps, removes or adds tokens; aux denotes any additional needed information. Such arbitrary attacks are expensive to optimize, and difficult to reason about. In practice, $T$ optimizes an adversarial suffix $x_{s}$ that is appended to $x_{p}$ to elicit $c$ Zou et al. (2023). These suffix-only attacks also allow efficient use of the KV-cache Pope et al. (2023). Specifically, we optimize the probability:
$Prob=P(c|x_{p}\oplus x_{s}).$ (1)
An adversary with white-box access to the LLM, can instead mount the attack in the embedding space i.e. modify the soft tokens:
$Prob=P(c|embed(x_{p})\oplus embed(x_{s})).$ (2)
In this case, $T$ uses the gradient from the LLM to update $x_{s}$. We visualize such attack in Figure 1.
Machine unlearning (MU) aims to remove information from models. Consider a machine learning model $f$ trained using a training dataset $D_{train}$. During an unlearning request to remove a specified subset $D_{forget}\in D_{train}$, the objective of MU is to produce an unlearned model $f_{u}$ that eliminates the influence of $D_{forget}$. There are two types of MU – exact, and approximate unlearning.
Exact unlearning ensures the output distribution of $f_{u}$ is statistically indistinguishable from $f_{ret}$ – a model retrained exclusively on the retained dataset $D_{retain}=D_{train}/D_{forget}$. This guarantees provable data removal, satisfying:
$p(f_{u}(x)=y)$ $=p(f_{ret}(x)=y)\quad$ (3)
$s.t.\quad\forall(x,y)\in D_{train}.$
It can be made more efficient by splitting the $D_{train}$ into overlapping chunks, and training an ensemble Bourtoule et al. (2021b). During an unlearning request, only the models containing the requested records are retrained. For certain classes of models, exact unlearning without retraining is possible, e.g. ECO adapts the Cauwenberghs and Poggio (CP) algorithm for exact unlearning within LeNet Huang et al. , and MUSE relabels the target data to achieve unlearning for over-parameterized linear models Yang et al. (2024).
Approximate unlearning relaxes the strict equivalence requirement, it only requires that $f_{u}$ approximates $f_{ret}$ within some bounded error. It relies on empirical metrics or probabilistic frameworks. In LLMs, approximate unlearning is typically accomplished by overwriting the information in the model Eldan and Russinovich (2023a); Wang et al. (2024), guiding the model away from it Feng et al. (2024), or editing the weights and/or activations Liu et al. (2024); Bhaila et al. (2024); Li et al. (2024); Tamirisa et al. (2024); Huu-Tien et al. (2024); Ashuach et al. (2024); Meng et al. (2022a, b).
Figure 1: STA combines $x_{p}$ with the optimized $x_{s}$.
3 Related work
While advances have been made in developing machine unlearning algorithms for LLMs, rigorous methodologies for auditing the efficacy of unlearning remain understudied. Adversarial soft token attacks (STAs) Schwinn et al. (2024) and 5-shot in-context prompting Doshi and Stickland (2024) have been shown to recover unlearned knowledge in models. When model weights can be modified, techniques such as model quantization Zhang et al. (2024e) and retraining on a partially unlearned dataset  Łucki et al. (2024); Hu et al. (2024); Chen et al. (2025) have also proven effective in recalling forgotten information.  Lynch et al. (2024) examined eight methods for evaluating LLM unlearning techniques and found that their latent representations remained similar. News and book datasets are used to analyze unlearning algorithms from six different perspectives Shi et al. (2024). It was shown that fine-tuning on unrelated data can restore information unlearned from the LLM Qi et al. (2024), indicating the existing unlearning methods do not actually remove the information but learn a refusal filter instead. Several benchmarks have been developed to evaluate the existing unlearning algorithms. Besides, an unlearning benchmark was introduced based on fictitious author information Maini et al. (2024a). For real-world knowledge unlearning, Real-World Knowledge Unlearning (RWKU) used 200 famous people as unlearning targets Jin et al. (2024), while WDMP focused on unlearning hazardous knowledge in biosecurity, cybersecurity Li et al. (2024).
4 Auditing with Soft Token Attacks
An oracle auditor $A_{o}$ takes an unlearned model $f_{u}$ and the candidate sentences $x_{c}\in X_{c}$, and outputs a ground truth, binary decision $a=\{0,1\}$ indicating whether the given records was part of $D_{train}$ of:
$a=A_{o}(f_{u},X_{c}=D_{forget},\textsl{aux}).$ (4)
$A_{o}$ is unrealistic but it can be easily instantiated for exact unlearning where $A_{o}$ knows the training data associated with $f$: $aux=\{D_{retain}\}$.
A realistic unlearning auditor $A_{u}$ takes an $f_{u}$, and $D_{forget}$ and outputs a score $s=(0,1)$ indicating whether the records were in $D_{train}$:
$s=A_{u}(f_{u},X_{c}=D_{forget},\textsl{aux}=\emptyset)$ (5)
$A_{u}$ represents cases where users unlearn facts from models that they did not create, e.g. to prevent harmful outputs.
In this work, we instantiate the soft token attack auditor $A_{\textsl{STA}}$ based on the soft token attacks (STAs) against unlearning Schwinn et al. (2024); Zou et al. (2024). $A_{\textsl{STA}}$ compares the relative difficulty of eliciting $c$ for $f_{ft}$ and $f_{u}$. The unlearning procedure is effective if eliciting completions using $f_{u}$ is more difficult than $f_{ft}$.
$s=A_{\textsl{STA}}(f_{u},X_{c}=D_{forget},\textsl{aux}=\{f_{ft}\}).$ (6)
Crucially, Schwinn et al., 2024 optimized $x_{s}$ only w.r.t. the affirmative beginning of $x_{c}$ ($x_{p}=$Who is Harry Potter?’, $x_{c}=$Harry Potter is’). Contrary to that minimal setting, our strong auditor $A_{\textsl{STA}}$ optimizes $x_{s}$ w.r.t. the entire $x_{c}$. Our goal is to study the extreme scenario where the auditor expects the exact completions for, e.g. to check for copyrighted content, or personal information.
In the Appendix, Figure 3 gives an overview of the auditing procedure, and the difference between $A_{o}$ and $A_{\textsl{STA}}$; Table 2 summarizes the notation.
5 Evaluation
5.1 Experiment setup
Datasets. For evaluation, we use two popular benchmark datasets: 1) Who Is Harry Potter?Eldan and Russinovich (2023a)(WHP) that intends to remove formation about the world of Harry Potter. WHP does not publish a full dataset. Hence, we use the snippets from the associated Hugging Face page, and we augment them with 2020 $(x_{p}\rightarrow c)$ Harry Potter trivia pairs generated with Llama2. 2) TOFUMaini et al. (2024a)is a dataset of fictional writers, designed to be absent from the LLM’s training data. Note that models released after TOFU was published might contain its records. We use the provided $10\%$ forget to $90\%$ retain split Maini et al. (2024b).
Models. We use Llama-2-7b-chat-hf Touvron et al. (2023) (Llama2), and Llama-3-8b-instruct Meta (2024) (Llama3). We get the unlearned WHP model from Hugging Face Eldan and Russinovich (2023b) (Llama2-WHP).
Implementation. We implement STA using LLMart Cornelius et al. (2025) – a PyTorch-based library for crafting adversarial prompts. We use implementations of the unlearning methods from the TOFU Maini et al. (2024c), and NPO Zhang et al. (2024b) repositories. We benchmark the attack against seven different unlearning algorithms: gradient ascent (GA), gradient difference (GDF) Liu et al. (2022), refusal (IDK) Rafailov et al. (2024), knowledge distillation (KL) Hinton (2015), negative preference optimization (NPO) Zhang et al. (2024c), NPO-GDF, NPO-KL.
5.2 Auditing with attacks
Who Is Harry Potter?. To elicit completions, we initialize the soft tokens using randomly selected hard tokens, and append them to the prompt $x_{a}=embed(x_{p})\oplus embed(x_{s})$. We then train the soft prompt using AdamW Loshchilov and Hutter (2019) for up to $3000$ iterations; using $lr=0.005$, and $\beta{s}=(0.9,0.999)$. $x_{p}$ does not change, only the embedded suffix does. If the optimization fails, we double the number of soft tokens up to the maximum of $16$. We report the mean and standard deviation over five independent runs across all prompts. In Table 1(a) we report the average number of soft tokens needed to elicit a completion. WHP * denotes the unlearned model with different prompt templates.
We show that all target completions can be generated with $\approx 4-6$ added soft tokens. For all pairs of models, we conduct a t-test under the null hypothesis $\mathcal{H}_{0}$ of equivalent population distributions with $\alpha=0.05$. We use an unpaired Welch’s t-test since sample variances are not equal WELCH (1947). We cannot reject the hypothesis for any of the pairs i.e. $p>0.05$. In other words, for all models, no significant evidence that eliciting completions is more difficult.
[…]
Prompt Model
template Llama2-WHP Llama3
N/A N/A $5.61\pm 6.32$
WHP $4.63\pm 3.69$ N/A
[…]
WHP +chat $4.12\pm 5.53$ N/A
(a) WHP results with different prompt templates.
Unlearning method Model
Llama2 Llama3
$f_{\emptyset}$ (none) $3.07\pm 3.25$ $3.11\pm 3.15$
$f_{ft}$ (none) $2.95\pm 3.35$ $3.21\pm 3.19$
$f_{u-IDK}$ $40\pm 3.20 3.3$$33\pm 3.09$
$f_{u-GA}$ $3.34\pm 3.97$ $3.21\pm 3.87$
$f_{u-GDF}$ $3.06\pm 3.34$ $3.11\pm 3.40$
$f_{u-KL}$ $3.08\pm 3.31$ $3.12\pm 3.17$
$f_{u-NPO}$ $3.11\pm 3.27$ $3.12\pm 3.27$
$f_{u-NPO-GDF}$ $3.15\pm 3.24$ $3.16\pm 3.16$
$f_{u-NPO-KL}$ $3.23\pm 3.62$ $3.24\pm 3.57$
(b) TOFU results with different unlearning methods.
Table 1: Number of soft tokens needed to elicit a completion for a fixed number of iterations; averaged over all prompts in each set and over five runs per prompt. When increasing the maximum iterations to 10,00010{,}000, all completions can be elicited with $1\text{--}2$ soft tokens.
TOFU. We follow the same setup as for WHP. In Table 1(b), we report the number of soft tokens required to elicit the completions. $f_{\emptyset}$ refers to the unmodified baseline model, $f_{ft}$ are the models fine-tuned on TOFU, followed by the unlearned models.
For all methods, we can elicit the completions with $\approx 3$ soft tokens. Similarly to WHP, for all pairs of models (within the same architecture), we conduct Welch’s t-test. We cannot reject the hypothesis for any of the pairs i.e. $p>0.05$; $f_{u-IDK}$ vs $f_{ft}$ (for Llama2) gives the lowest p-value of $0.509$. For all models, there is not enough evidence to say that eliciting completions is more difficult.
One could argue that the unlearning methods used are not effective (when comparing $f_{ft}$ vs $f_{u-*}$), hence they require similar numbers of soft tokens. In fact, most of these approaches have already been shown to be ineffective and susceptible to simple paraphrasing Shi et al. (2024). However, the same holds when compared to $f_{\emptyset}$. In the next section, we demonstrate that the result cannot be attributed to the (in-)effectiveness of the unlearning methods but rather the power of STA.
5.3 Eliciting random strings
The chance of a random string appearing in the training set is negligible, and preceding tokens do not inform the selection of the next token. We construct random strings uniformly at random from the range 33-126 of the ASCII table.
We initialize the soft prompt using randomly selected tokens. In this experiment, there is only $x_{s}$, and no $x_{p}$. We then train the soft prompt using AdamW for up to $3000$ iterations per soft token; using $lr=0.005$, and $\beta{s}=(0.9,0.999)$.
In Figure 2, we report the longest elicited string for a given number of soft tokens. We repeat the experiment five times – e.g., the first marker implies that for each of the five tested random strings of length $150$, we found an effective soft prompt. We observe that not all initializations and seed configurations succeed, in which case a run needs to be restarted with a different seed. If the loss plateaus around $25\%$ of the iterations, we restart the run. However, no single string was restarted more than ten times. Our results show that STAs can be used to elicit completely random strings, thus undermining their application for auditing unlearning.
Next, we aim to answer why eliciting strings is possible. Prompt-tuning Lester et al. (2021) is an efficient fine-tuning technique that trains only a soft prompt added to the input instead of all weights. STAs can be viewed as an extreme case of prompt-tuning, where instead of training over many prompts, one trains an attack per each prompt. Thus, an LLM that outputs a completion that it was trained on is an expected behavior. However, one could argue, in practice, a properly unlearned (or aligned) LLM should never output undesirable text.
Figure 2: A single soft token can force over $150$ random characters. With $10$ soft tokens, it is possible to generate over $400$ random characters.
6 Discussion & Conclusion
Unlearning vs jailbreaking. Our findings also apply to the jailbreaking community. Prior work hinted that unlearning and preventing harmful outputs can be both viewed as suppressing particular information Zhang et al. (2024d). For instance, it was shown that fine-tuning on benign or unrelated data can restore undesirable behavior Hu et al. (2024); Łucki et al. (2024).
Variation in gradient descent. Prior work showed that retraining with some records removed can result in the same final model depending on the seed Thudi et al. (2022). The influence of the records might be minimal, making unlearning unnecessary. Similarly, it was shown that SGD has intrinsic privacy guarantees, assuming there exists a group of similar records Hyland and Tople (2022). Thus, algorithmic auditing of unlearning might not be possible, and one would have to rely on verified or attested procedures instead Eisenhofer et al. (2023), regardless of their impact on the model.
Distinguishing learned soft tokens. Even though, in most of our results, the number of soft tokens required to elicit a completion is the same, we attempt to distinguish between them. To this end, we take all single-token STAs optimized for TOFU (Table 1(b)) and assign a label $y=\{0,1\}$: $y=0$ for $f_{\emptyset}$, and $y=1$ for $f_{ft}$ and the unlearned models. We then train a binary classifier using $f_{\emptyset}$ and $f_{ft}$. While we are able to overfit it and distinguish between $f_{\emptyset}$ and $f_{ft}$, we were not able to train a model that would generalize to the unlearned models, and decisively assign a class. Our approach is similar to Dataset Inference Maini et al. (2021, 2024d) which showed there can be distributional differences between the models, depending on the data they were trained on. Further investigation into what soft tokens are learned during the audit is an interesting direction for future work.
Conclusion. In this work, we show that soft token attacks (STA) cannot distinguish between base, fine-tuned, and unlearned models: a strong auditor can elicit all unlearned text. Also, we show that STA with a single token can elicit $150$ random characters, and over $400$ with $10$ tokens. Our work shows that machine unlearning in LLMs needs careful assumptions to avoid misleading results.
7 Limitations & ethical considerations
Limitations. Our experiments are constrained to models with 7–8 billion parameters. Nevertheless, since the expressive power of LLMs generally scales with size Kaplan et al. (2020), we expect our findings to extend to larger models. For efficiency reasons, we restrict our evaluation to at most $10$ soft tokens and random strings of up to $400$ characters, which does not establish an upper bound on the length of strings that can be elicited. Future work could investigate whether black-box optimization methods—such as zeroth-order optimization Chen et al. (2017)—can reproduce the elicitation observed in the white-box setting. In addition, extending our evaluation with random strings may help determine whether there exists a clear and generalizable relationship between the number of soft tokens and the maximum number of generated characters.
Ethical considerations. In this work, we show that an auditor with white-box access and sufficient computational resources can elicit arbitrary text from an LLM. Although this requires knowledge of the target completion, partial completions may suffice, enabling the extraction of harmful information—especially when the auditor has approximate prior knowledge of the removed content.
8 Acknowledgments
We thank Cory Cornelius and Marius Arvinte for valuable discussions on this work.
Appendix A Auditing process
Figure 3 gives a complete overview of the auditing procedure, and the difference between $A_{o}$ and $A_{\textsl{STA}}$. In Table 2, we summarize the notation.